OpenStudy (anonymous):

Solve K = (b – 2a) for b

OpenStudy (diyadiya):

Just add 2a on both sides

OpenStudy (diyadiya):

what do you get now?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

b=K+2a

OpenStudy (diyadiya):

K = b-2a we have to isolate b so we add 2a on both sides K + 2a = b - 2a + 2a *** -2a + 2a = 0*** K+2a = b